OpenStudy (anonymous):

FAN AND MEDAL!!!!!! Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (perl):

hints: lines that are parallel have the same slope

OpenStudy (anonymous):

I know that

OpenStudy (perl):

ok so you know that your line is y = -3x + b , now you need to find b

OpenStudy (anonymous):

how would I do that?

OpenStudy (perl):

you can plug in the given point (2,6) for x and y, and solve for b

OpenStudy (anonymous):

6= -3(2) -4?

OpenStudy (anonymous):

hello?

OpenStudy (perl):

it should be

OpenStudy (perl):

y = -3x + b , plug in (2,6) 6 = -3*2 + b 6 = -6 + b 6 + 6 = b 12 = b